How would you find x and y by elimination of substituting this problem?
x+2y=-5
4x-3y=2

Answer by sdmmadam@yahoo.com(530)

How would you find x and y by elimination of substituting this problem?
x+2y=-5 ----(1)
This gives x = (-5-2y) ----(*) (adding (-2y) to both the sides)
Putting x= (-5-2y) in (2)
4x-3y=2 ----(2)
4X(-5-2y)-3y = 2
-20-8y-3y = 2
-20-11y = 2
-20-2 = 11y (adding 11y and -2 to both the sides)
-22 = 11y
y = -22/11= -2
y =-2 in (*) gives x= -5-2y =-5-2X(-2) = -5+4 =-1
Answer: x= -1 and y = -2
Verification: x= -1 and y = -2 in (1)
Then LHS = x+2y = -1+2X(-2) =-1-4 = -5=RHS